OpenStudy (anonymous):

Hi, for work - energy theorem, is it just; W = 1/2mv _{f}^2-1/2mv _{0}^2 or is there many ways of it like is this also part of work energy theorem? W = mg(h _{0}-h _{f})=1/2mv _{0}^2

OpenStudy (anonymous):

\[W = 1/2mv _{f}^2-1/2mv _{0}^2\]

OpenStudy (anonymous):

\[W = mg(h _{0}-h _{f})=1/2mv _{0}^2\]

OpenStudy (anonymous):

Long as any W = any kind of energy would that be work energy theorem.

OpenStudy (anonymous):

Or is the work energy theorem only \[W =KE _{f}-KE _{0}\]

OpenStudy (anonymous):

@wio Mind answering this?

OpenStudy (anonymous):

Work energy theorem is just: \[ W_{net}=\Delta E \]

OpenStudy (anonymous):

Ah that's what I thought. Thanks haha.